OpenStudy (richgall____):

Can someone help with this last question 😊A triangle has sides 8 centimeters, 10 centimeters, and 15 centimeters. Is the triangle a right triangle?

OpenStudy (richgall____):

a^2+b^2=c^2 yea ik

OpenStudy (richgall____):

8^2=64 10^2=100 15^2=225

OpenStudy (mathmale):

We need to determine whether or not this is a right triangle. Immediately look at the longest side. If this triangle is a right triangle, then the longest side is also the hypotenuse. Pythagorean Theorem states that if you square the lengths of both shorter sides, and then sum up these squares, that sum equals the square of the longest side. So, either the Pythagorean Theorem equation will be true or it will be false. If false, the triangle is not a right triangle. If true, the triangle is a right triange.

OpenStudy (mathmale):

Please show your work.

OpenStudy (richgall____):

i already did in the recent post

OpenStudy (richgall____):

8^2+10^2=15^2

OpenStudy (mathmale):

Good. Now determine whether or not that equation is true.

OpenStudy (richgall____):

it's not one because 64+100 is not 15^2

OpenStudy (mathmale):

Very good. That's correct.

OpenStudy (richgall____):

thanks.

OpenStudy (mathmale):

happy to be of help. Good going.